OpenStudy (anonymous):

2,6,14,24,42,62,? complete the sequence and also tell how it is solved.

OpenStudy (anonymous):

Is the 4th number supposed to be 26?

OpenStudy (anonymous):

because then you'd have changes of +4, +8, +12, +16, +20, which makes sense. The next number would be 62+24=86. As you wrote it, it doesn't really make sense.

jhonyy9 (jhonyy9):

pardon but not is correct how you have wrote it because 14+12 not equal 24 - is 26

jhonyy9 (jhonyy9):

- so but i think that is : 2+4=6 6+4+4=14 14+6+4=24 24+14+4=42 42+24-4=62 62+42-4=100 100+62-4=158 158+100-4=254